Use the quadratic formula to solve x² + 7x + 8 = 0.
What are the solutions to the equation?
Round irrational solutions to the nearest tenth.

Answer by stanbon(75887)

Use the quadratic formula to solve x² + 7x + 8 = 0.
What are the solutions to the equation?
Round irrational solutions to the nearest tenth.
------
x = [-7+-sqrt(49-4*8)]/(2)
----
x = [-7+-sqrt(17)]/2
-----
x = (-7+sqrt(17))/2 or x = (-7-sqrt(17))/2
----
x = (-7+4.1)/2 or x = (-7-4.1)/2
-----
x = -1.45 or x = -5.55
----------
Cheers,
Stan H.
----------